Two free positive charge 4q and quite are a distance l apart. What charge Q is needed to achieve equilibrium for the entire system and where should it be placed from charge q

Grade:12

1 Answers

Smartkids Jind
6 years ago
We have 2same charges(+ve) so their force will balance on the line joining.4q_______________q _________ ______ l-x. x. Let the neutral position is x distance from qAt x distance Q is placed .Force here is balanced:k4qQ/(l-x)²=kqQ/x²x=l/3 kkqQ/(l/3)²=k4q.q/l²From this equation we will get Q